OpenStudy (anonymous):

solve the equation and check your solution 15n^2-n=2

OpenStudy (anonymous):

25n^2-n-2 = 0 (rearrange terms) (3n + 1)(5n - 2)=0 Solve for both n = -1/3, 2/5 (these are the roots)

OpenStudy (anonymous):

Er, 15n^2-n-2 = 0 for top line